OpenStudy (chasebraves):

Given: F(x) = x + 2 and G(x) = 3x + 5 (F + G) (x) = 3x + 7 4x + 7 3x² + 7

TheSmartOne (thesmartone):

(f + g)(x) = f(x) + g(x) = x + 2 + 3x + 5 = ??

OpenStudy (chasebraves):

4x + 7

OpenStudy (chasebraves):

@TheSmartOne is it tight

TheSmartOne (thesmartone):

That is correct! Great job! :)